The amount of time between taking pain reliever and getting relief is normally distibuted with a mean of 23 & a standard deviation of 4 minutes. Find the probability that the time between taking the medication and getting relief is at least 17 minutes
Answer #1

Here, μ = 23, σ = 4 and x = 17. We need to compute P(X >= 17). The corresponding z-value is calculated using Central Limit Theorem

z = (x - μ)/σ
z = (17 - 23)/4 = -1.5

Therefore,
P(X >= 17) = P(z <= (17 - 23)/4)
= P(z >= -1.5)
= 1 - 0.0668
= 0.9332